What is maximum height expression for projectile motion.
1
Expert's answer
2020-10-07T10:16:13-0400

The maximum height for projectile motion:


"h_{max}=\frac{v_0^2sin^2\alpha}{2g}"


where "v_0" - the initial velocity,

"\alpha" - the angle at which the projectile is moved above the horizon.

"g=10\frac{m}{s^2}"